what mass of chromium metal can be plated on the cathode from a CᵣCₗ₃ solution using a current of 5.8A for 2 minutes?
1
Expert's answer
2020-01-06T04:50:38-0500

Cathode: "Cr^{3+} + 3e^- \rightarrow Cr^0"


"m(Cr) = Z\times I \times t = \frac{52.00 g Cr}{3(96500)} \times \frac{5.8 C}{s} \times 2 min \times \frac{60 s}{1 min} = 0.125 g"